# Small Gradient Norm Regret for Online Convex Optimization

### v1

@@ 2. A2:
A1 is slightly different from the standard smoothness assumption $\ell_{t}(x)-\ell_{t}(y)-\langle\nabla\ell_{t}(y),x-y\rangle\leq\frac{L}{2}\|x-y\|^{2}$: a smooth convex function on $\mathbb{R}^{n}$ is $\frac{1}{L}$-strongly convex in the dual space. It yields a tighter lower bound than that provided by convexity alone [25]. This sharper lower bound is central to the derivation of the $G^{\star}$ regret.
@@ 2.2 Upper and lower bound for the $G^{\star}$ regret
This section presents algorithms that achieve an $\mathcal{O}(\sqrt{G^{\star}})$ regret upper bound with a matching lower bound. The key idea is simple. Consider the regret analysis of a prototypical online learning algorithm, such as online gradient descent (OGD) with constant learning rate $\eta>0$:
@@
where (6) uses non-expansiveness of projection, (7) applies convexity $\ell_{t}(x)\geq\ell_{t}(x^{t})+\langle\nabla\ell_{t}(x^{t}),x-x^{t}\rangle$, and the $\|\nabla\ell_{t}(x^{t})\|^{2}$ term is further upperbounded in (8) using self-boundeness of (non-negative) $L$-smooth functions. However, a smooth function over $\mathbb{R}^{n}$ (A1) satisfies as a stronger inequality
@@ 3.2 $G^{\star}$ regret under bandit feedback
Under A1, A2 and suppose $n\geq 8$, bandit gradient descent (Algorithm˜5 in the appendix) $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta g_{t}]$ with constant learning rate $\eta\in(0,\tfrac{1}{4nL})$ satisfies $\mathbb{E}[\rho_{T}(x)]\leq\textstyle\tfrac{D^{2}}{2\eta}+\tfrac{4n\eta}{1-4n\eta L}\sum_{t=1}^{T}\|\nabla\ell_{t}(x)\|^{2}+\tfrac{4n\eta}{1-4n\eta L}TL^{2}\mu^{2}+\tfrac{\eta T}{2}n^{2}L^{2}\mu^{2}+\tfrac{TL\mu^{2}}{2}$ for any $x\in\mathcal{X}$. In particular, choosing $x=x^{\star}$, $\eta=\min\big\{\tfrac{D}{4\sqrt{nG^{\star}_{T}}},\tfrac{1}{8nL}\big\}$ and $\mu=\tfrac{D}{\sqrt{2nT}L}$ gives $\mathbb{E}[\rho_{T}(x^{\star})]\leq\max\{8nLD^{2},4D\sqrt{nG^{\star}_{T}}\}+\tfrac{LD^{2}}{2}$.
@@
Under the same assumptions as Theorem˜3.3, bandit gradient descent (Algorithm˜6 in the appendix) $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta_{t}g_{t}]$ with AdaGrad-Norm learning rate $\eta_{t}=\tfrac{\alpha}{\sqrt{\sum_{i=1}^{t}\|g_{i}\|^{2}}}$, $\alpha=\tfrac{\sqrt{2}D}{2}$ and $\mu=\tfrac{D}{n\sqrt{T}}$ satisfies $\sum_{t=1}^{T}\mathbb{E}[\rho_{T}(x)]\leq\max\{16LD^{2}n,4D\sqrt{nG_{T}(x)}\}+5LD^{2}$ for any $x\in\mathcal{X}$.
@@ 1. A3:
    $\mathbb{E}[f(\bar{x}_{T})-f(x^{\star})]\leq\tfrac{1}{2\eta T}\|x^{1}-x^{\star}\|^{2}+\tfrac{\eta}{2(1-\eta L)}(\sigma_{g}^{\star})^{2};$
  • if AdaGrad-Norm from Theorem˜2.2 is applied to $\{\ell_{t}\}_{t=1}^{T}$, then $\mathbb{E}[f(\bar{x}_{T})-f(x^{\star})]\leq\tfrac{LD^{2}}{T}+\tfrac{\sqrt{2}D}{\sqrt{T}}\sigma_{g}^{\star}$.
  • if AdaFTRL from Theorem˜2.3 is applied to $\{\ell_{t}\}_{t=1}^{T}$, then $\mathbb{E}[f(\bar{x}_{T})-f(x^{\star})]\leq\tfrac{2R^{2}LD^{2}}{T}+\tfrac{\sqrt{3}RD}{\sqrt{T}}\sigma_{g}^{\star}$.
where $\bar{x}_{T}\coloneqq\tfrac{1}{T}\sum_{t=1}^{T}x^{t}$ is the average iterate and $R$ is defined in Theorem˜2.3.
@@
Suppose AdaGrad-Norm is applied to (15) with $\mathcal{X}=\mathbb{R}^{n}$ and that there exists $x^{\star}$ such that $\nabla f(x^{\star},\xi)=0$ surely. Further assume that each $f(x,\xi)$ is $L$-smooth. Then $\mathbb{E}[f(\bar{x}^{T+1})-f(x^{\star})]\leq\tfrac{L}{2T}(\tfrac{\|x^{1}-x^{\star}\|^{2}}{2\alpha}+\alpha\log(\tfrac{\sqrt{\mathrm{e}}\alpha L}{\|\nabla f(x^{1},\xi^{1})\|})+\tfrac{1}{2L}\|\nabla f(x^{1},\xi^{1})\|)^{2}$, where $\bar{x}_{T}\coloneqq\tfrac{1}{T}\sum_{t=1}^{T}x^{t}$ is the average iterate.

### v2

@@ 2. A2:
A1 looks slightly different from, but is equivalent to, the standard smoothness assumption $\ell_{t}(x)-\ell_{t}(y)-\langle\nabla\ell_{t}(y),x-y\rangle\leq\frac{L}{2}\|x-y\|^{2}$: an $L$-smooth convex function on $\mathbb{R}^{n}$ is $\frac{1}{L}$-strongly convex in the dual space. It yields a tighter lower bound than that provided by convexity alone [25]. This sharper lower bound is central to the derivation of the $G^{\star}$ regret.
@@ 2.2 Upper and lower bound for the $G^{\star}$ regret
This section presents algorithms that achieve an $\mathcal{O}(\sqrt{G^{\star}_{T}})$ regret upper bound with a matching lower bound. The key idea is simple. Consider the regret analysis of a prototypical online learning algorithm, such as online gradient descent (OGD) with constant learning rate $\eta>0$:
@@
where (6) uses non-expansiveness of projection, (7) applies convexity $\ell_{t}(x)\geq\ell_{t}(x^{t})+\langle\nabla\ell_{t}(x^{t}),x-x^{t}\rangle$, and the $\|\nabla\ell_{t}(x^{t})\|^{2}$ term is further upperbounded in (8) using self-boundeness of (non-negative) $L$-smooth functions. However, a smooth function over $\mathbb{R}^{n}$ (A1) satisfies a stronger inequality
@@ 3.2 $G^{\star}$ regret under bandit feedback
Under A1, A2 and suppose $n\geq 8$, bandit gradient descent (Algorithm˜5 in the appendix) $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta g_{t}]$ with constant learning rate $\eta\in(0,\tfrac{1}{4nL})$ satisfies $\mathbb{E}[\rho_{T}(x)]\leq\textstyle\tfrac{D^{2}}{2\eta}+\tfrac{4n\eta}{1-4n\eta L}G_{T}(x)+\tfrac{4n\eta}{1-4n\eta L}TL^{2}\mu^{2}+\tfrac{\eta T}{2}n^{2}L^{2}\mu^{2}+\tfrac{TL\mu^{2}}{2}$ for any $x\in\mathcal{X}$. In particular, choosing $x=x^{\star}$, $\eta=\min\big\{\tfrac{D}{4\sqrt{nG^{\star}_{T}}},\tfrac{1}{8nL}\big\}$ and $\mu=\tfrac{D}{\sqrt{2nT}L}$ gives $\mathbb{E}[\rho_{T}(x^{\star})]\leq\max\{8nLD^{2},4D\sqrt{nG^{\star}_{T}}\}+\tfrac{LD^{2}}{2}$.
@@
Under the same assumptions as Theorem˜3.3, bandit gradient descent (Algorithm˜6 in the appendix) $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta_{t}g_{t}]$ with AdaGrad-Norm learning rate $\eta_{t}=\tfrac{\alpha}{\sqrt{\sum_{i=1}^{t}\|g_{i}\|^{2}}}$, $\alpha=\tfrac{\sqrt{2}D}{2}$ and $\mu=\tfrac{D}{n\sqrt{T}}$ satisfies $\mathbb{E}[\rho_{T}(x)]\leq\max\{16LD^{2}n,4D\sqrt{nG_{T}(x)}\}+5LD^{2}$ for any $x\in\mathcal{X}$.
@@ 1. A3:
    $\mathbb{E}[f(\bar{x}^{T})-f(x^{\star})]\leq\tfrac{1}{2\eta T}\|x^{1}-x^{\star}\|^{2}+\tfrac{\eta}{2(1-\eta L)}(\sigma_{g}^{\star})^{2};$
  • if AdaGrad-Norm from Theorem˜2.2 is applied to $\{\ell_{t}\}_{t=1}^{T}$, then $\mathbb{E}[f(\bar{x}^{T})-f(x^{\star})]\leq\tfrac{LD^{2}}{T}+\tfrac{\sqrt{2}D}{\sqrt{T}}\sigma_{g}^{\star}$.
  • if AdaFTRL from Theorem˜2.3 is applied to $\{\ell_{t}\}_{t=1}^{T}$, then $\mathbb{E}[f(\bar{x}^{T})-f(x^{\star})]\leq\tfrac{2R^{2}LD^{2}}{T}+\tfrac{\sqrt{3}RD}{\sqrt{T}}\sigma_{g}^{\star}$.
where $\bar{x}^{T}\coloneqq\tfrac{1}{T}\sum_{t=1}^{T}x^{t}$ is the average iterate and $R$ is defined in Theorem˜2.3.
@@
Suppose AdaGrad-Norm is applied to (15) with $\mathcal{X}=\mathbb{R}^{n}$ and that there exists $x^{\star}$ such that $\nabla f(x^{\star},\xi)=0$ surely. Further assume that each $f(x,\xi)$ is $L$-smooth. Then $\mathbb{E}[f(\bar{x}^{T})-f(x^{\star})]\leq\tfrac{L}{2T}(\tfrac{\|x^{1}-x^{\star}\|^{2}}{2\alpha}+\alpha\log(\tfrac{\sqrt{\mathrm{e}}\alpha L}{\|\nabla f(x^{1},\xi^{1})\|})+\tfrac{1}{2L}\|\nabla f(x^{1},\xi^{1})\|)^{2}$, where $\bar{x}^{T}\coloneqq\tfrac{1}{T}\sum_{t=1}^{T}x^{t}$ is the average iterate.